OpenStudy (abbycross167):

Can someone please check my work? Please? I would really appreciate it... Question~ Solve the following problem using any method of solving systems. CHECK your solutions by substituting your coordinates into BOTH equations and simplifying. In one week a music store sold 7 violins for a total of $1600. Two different types of violins were sold. One type cost $200 and the other type cost $300. How many of each type of violin did the store sell?

OpenStudy (abbycross167):

My answer~ 4 violins were sold at $300 = $1200 2 were sold at $200 = $400 $1200+$400= $1600

OpenStudy (abbycross167):

What my teacher said was wrong~ This is not a guess and check problem. REview the real-life example at the end of the lesson and use it as a model for this.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it is wrong because 4 violins plus 2 violins is 6 violins and they sold 7

OpenStudy (abbycross167):

If you would like to see the "real life example" that my teacher is talking about, I will be happy to post that! And hello @misty1212

OpenStudy (misty1212):

you want to go ahead and solve it? we need a couple of equations

OpenStudy (misty1212):

lol no i really don't want to solve this one ? (none of these has anything to do with "real life')

OpenStudy (abbycross167):

I'm not too sure how to...

OpenStudy (misty1212):

we need a variable or two variables depending on your teacher

OpenStudy (misty1212):

we could say \(x\) is the number of $300 violins and y is the number of $200 violins and then make two equations

OpenStudy (misty1212):

since the total number of violins sold is 7, we know \[x+y=7\]

OpenStudy (misty1212):

and the total money is $1600 so \[300x+200y=1600\]

OpenStudy (abbycross167):

oh ok

OpenStudy (misty1212):

the first equation is the same as \[y=7-x\] so put it in the second and solve \[300x+200(7-x)=1600\]

OpenStudy (abbycross167):

ok, could you show me how?

OpenStudy (abbycross167):

@misty1212

OpenStudy (abbycross167):

x=2?

OpenStudy (mathstudent55):

300x + 1400x - 200x = 1600 100x = 200 x = 2 Correct. Now substitute x = 2 in the equation x + y = 7 and find y.

OpenStudy (abbycross167):

so 2x+y=7? or y+2x=7?

OpenStudy (mathstudent55):

Neither. Since x = 2, replace x by 2. x + y = 7 Rewrite the equation above, substituting 2 for x. Then solve for y.

OpenStudy (abbycross167):

2+y=7= 5 y=5?

OpenStudy (abbycross167):

@mathstudent55 ?

OpenStudy (mathstudent55):

You need to do one step at a time and write always true statements. The equation we start with is x + y = 7 We know x = 2, so we replace x with 2: 2 + y = 7 Now we subtract 2 from both sides to isolate y: 2 + y - 2 = 7 - 2 y = 5 You are correct. y = 5, but you cannot write: 2+y=7= 5 because 7 is not equal to 5.

OpenStudy (mathstudent55):

Since we get x = 2 and y = 5, now we see what we had these two variables represent, so we can answer the question.

OpenStudy (mathstudent55):

We look at @misty1212 's statement above: "we could say x is the number of $300 violins and y is the number of $200 violins and then make two equations " Since we got x = 2 and y = 5, now we answer: The store sold 2 $300 violins and 5 $200 violins Let's check: 2 + 5 = 7, so the number fo violins sold is correct. 2 * $300 + 5 * $200 = $600 + $1000 = $1600 The amount of money also checks, so our solution is correct.

OpenStudy (abbycross167):

oh ok thank you soo much!! you are a great help!

OpenStudy (mathstudent55):

You're welcome.